Project

General

Profile

Nightly Observing Plan for MJD 57400 (starting evening of 12/01/2016)

DES (2012B-0001)

Observers

  • Anthony Kremin (Observer 1)
  • Vinu Vikraman (Observer 2/RM)

Shift list

CTIO Staff

  • Manuel Hernández

Staff calendar

Almanac

mjd 57400, date 20160112, Site CTIO, UT-offset 3 hr
Current local time 14:43 CLST     UT 17:43     SidTim 20:28
Sunset     20:56 CLST  23:56 UT   SidTim   40 deg  
Eve. -10   21:36 CLST  00:36 UT   SidTim   50 deg  Zeropoint; standards
Eve. -14   21:59 CLST  00:59 UT   SidTim   56 deg  Begin obstac
Eve. -18   22:22 CLST  01:22 UT   SidTim   62 deg  Last chance standards
Moonset    22:45 CLST  01:45 UT   SidTim   67 deg  
Midpoint   01:51 CLST  04:51 UT   SidTim  114 deg  
Morn -18   05:20 CLST  08:20 UT   SidTim  166 deg  Begin astronomical twilight
Morn Std   05:44 CLST  08:44 UT   SidTim  172 deg  End obstac; begin standards
Morn -10   06:06 CLST  09:06 UT   SidTim  178 deg  
Sunrise    06:46 CLST  09:46 UT   SidTim  188 deg  
Moon position (deg): RA 333.7, Dec -8.2 Phase: 0.203

Weather Conditions

Observing Plan and Goals for the Night

  • evening standards (if clear) [Yzirg ordered scripts]
UT Date = '2016/01/13 00:36'     Local time = 21:36
SDSSJ0000-0000        ra 00:00:00  dec +00:00:00  secz 1.8
SDSSJ0100-0000        ra 01:00:00  dec +00:00:00  secz 1.4
SDSSJ0320-0000        ra 03:20:00  dec +00:00:00  secz 1.2
  • Obstac
  • No SN on deadman but many that could be triggered on seeing.
  • Middle of the night standards (if clear) [grizY ordered scripts]
UT Date = '2016/01/13 04:31'     Local time = 01:31
SDSSJ1048-0000        ra 10:48:00  dec +00:00:00  secz 1.9
SDSSJ0843-0000        ra 08:43:00  dec +00:00:00  secz 1.2

Notes from the 4PM Meeting and Additional Comments.

Camera & Telescope
No problems. Continuation of the dome bearing work. A little trouble, not usual, with TCS at the moment. Should be OK in a few minutes.

Last Night: http://des-ops.fnal.gov:8080/nightsum/nightsum-2016-01-11/nightsum.html
Nice at the beginning w/ 1" FWHM. Then degrading to 1.3", humidity went high. C3g SN was observed, seeing triggered.

DESDM has about 30 exposures. ~84 still missing. DTS is investigating. N word on the dq.

Notes from Observers

See Night summary

Notes on Analyses of Night's Data: